OpenStudy (anonymous):

Consider the graph of y = ( x + 2)^2 - 4. Find the inverse of the equation.

OpenStudy (rogue):

\[y = ( x + 2)^2 - 4\]To find the inverse, switch the variables, and solve for y.\[x = ( y + 2)^2 - 4\]\[x + 4 = (y + 2)^2\]\[\pm \sqrt {x + 4} = y + 2\]\[y = -2 \pm \sqrt {x + 4}\]